A circle is inscribed in a square of side 14cm.What is the ratio of the area of the circle and the square?Please explain the answer in details.

Grade:7

3 Answers

Vijay Mukati
8 years ago
Dear Student.

Given the side length of square is 14cm, which means the radius of inscribed circle will be half i.e. 7 cm.
Now the area of Cirlce is (22/7)*7*7 = 154cm2.
and the area of square is 14*14 cm2

Therefore its ratio will be = 154/(14*14).

Just simpilify it now.

Thanks.
Abhishek Chaudhary
8 years ago
Area of square=14*14
Area of circle=7*7*22/7=22*7
Ratio=22*7/14*14
=22/14*2
=11/14
Ratio is equal to 11/14
vijay kale
6 years ago
The circle is inscribed inside the square means that the circle is placed inside the square in such a manner that it touches every side of the square. So the diameter of the circle is equal to side of the square.Therefore radius of circle is 7cm which is half of 14cm.Area of square=side^2=14*14Area of circle=pie×r×r=7*7*22/7=22*7Ratio=22*7/14*14=22/14*2=11/14Required Ratio is 11/14